# Fast Equivariant Imaging: Acceleration for Unsupervised Learning via Augmented Lagrangian and Auxiliary PnP Denoisers

### v1

@@ Abstract
We propose Fast Equivariant Imaging (FEI), a novel unsupervised learning framework to efficiently train deep imaging networks without ground-truth data. From the perspective of reformulating the Equivariant Imaging based optimization problem via the method of Lagrange multipliers and utilizing plug-and-play denoisers, this novel unsupervised scheme shows superior efficiency and performance compared to vanilla Equivariant Imaging paradigm. In particular, our PnP-FEI scheme achieves an order-of-magnitude (10x) acceleration over standard EI on training U-Net with CT100 dataset for X-ray CT reconstruction, with improved generalization performance.
@@ 1 Introduction
$y=Ax^{\dagger}+\epsilon,$ (1)
with reference signal / image $x^{\dagger}\in\mathbb{R}^{n}$ to be estimated, noisy measurements $y\in\mathbb{R}^{m}$, measurement noise $\epsilon\in\mathbb{R}^{m}$. $A\in\mathbb{R}^{m\times n}$ is the forward operator potentially with $m\ll n$, and can be a downsampling operator for super-resolution task, masks for inpainting, Gaussian linear measurements for compressive sensing, Radon transform for CT/PET, to name a few. In this setting, one can construct an MAP estimator for $x^{\dagger}$ by solving the following optimization problem with classical proximal splitting algorithms (Combettes and Pesquet, 2011):
@@
where $f_{\mathrm{mc}}(x)$ is the measurement consistency (MC) term such as $\ell_{1}$ loss, $\ell_{2}$ loss and SURE-loss (Chen et al., 2022), $R(x)$ is the prior or regularizer.
In unsupervised learning, given that we only have access to noisy measurements $\mathcal{Y}$, one can reconstruct the signal/image by learning a reconstruction function $G_{\theta}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ such that $G_{\theta}(y)\approx x^{\dagger}$. By enforcing that the inverse mapping $G_{\theta}$ is consistent in the measurement domain, the following measurement consistence estimator (MC) can be obtained,
@@
However, the MC constraint is not enough to learn the inverse mapping, as it cannot capture information about the signal/image outside the range of the operator $A^{\top}$, we refer the interested reader to Chen et al. (2021) for details.
In order to learn beyond the range space of $A^{\top}$, Chen et al. (2021) takes advantage of some mild prior information mainly about the symmetries of plausible signal/image, $i.e.$, signal/image is invariant to certain groups of transformations, such as shifts, rotations, reflections, etc. According to invariance, the composition $G_{\theta}\circ A$ should be equivariant to the transformation $T_{g}$, $i.e.$
@@
This deep-learning-based unsupervised method has recently achieved state-of-the-art performance for a host of inverse problems (Chen et al., 2021, 2022; Terris et al., 2024). Their success has been attributed to their ability to explore group symmetry from a range of physics, at the cost of cumbersome and time-consuming training. However, one of the main challenges of this type of unsupervised schemes is the computational cost of training due to the slow convergence and large computational complexity across training iterations, particularly for large-scale inverse problems and advanced architectures such as unrolling (Monga et al., 2021; Zhou et al., 2024). In our recent work of Sketched EI (Xu et al., 2024) we partially addressed the computational complexity issue for those inverse problems on which dimensionality reduction strategies can be applied. In this work we aimed to address the problem of slow convergence of standard optimizers such as Adam (Kingma and Ba, 2015) on minimizing this EI objective.
In this work, we explore from an optimization perspective to accelerate this promising unsupervised method via proximal alternating minimization (Afonso et al., 2010), namely Fast Equivariant Imaging (FEI), for faster convergence. Our modification is achieved using an Lagrangian formulation of the problem, and is solved using an alternating-direction minimization method. From numerical experiments, we show that this fast EI method and its PnP variant (PnP-FEI) can improve both the convergence rate and performance.
@@ 1.1 Relation to Prior Works
We differ from Chen et al. (2021) in two important ways: First, the way they minimize the objective function is by optimizing the data consistency term and equivariance term uniformly in a single backpropagation, which hence enjoys moderate convergence and time-consuming training. In contrast, we consider a more aggressive minimization for the data consistency term via proximal alternating minimization. A second crucial difference is that within our FEI scheme, we can easily incorporate plug-and-play framework as implicit priors to further speed the convergence. In sharp contrast, the vanilla EI, by only exploiting group symmetry of the physics, can be seen as a special case of our proposed PnP-FEI. Although we focus only on accelerating EI in this work, the same algorithmic structure can be easily extended to more recent variants of EI such as Robust Equivariant Imaging (REI) (Chen et al., 2022), Multi-Operator Imaging (MOI) (Tachella et al., 2022), and our recently proposed Sketched Equivariant Imaging (SkEI) (Xu et al., 2024) which utilizes the dimensionality reduction techniques for acceleration.
@@ 2 Proposed Methods
2.1 Fast Equivariant Imaging
We propose solving the problem (4) from an optimization point of view, which leads us to reformulate (4) as follows
@@ 2.1 Fast Equivariant Imaging
Typical choices for the measurement consistency term $f_{\mathrm{mc}}(x)$ includes $\ell_{1}$ loss, $\ell_{2}$ loss or SURE-loss (Chen et al., 2022), etc. We simply substitute $G_{\theta}(y)$ with $\mathbf{u}$ in the MC term and further incorporate the EI regularizer. From another perspective, the neural network $G_{\theta}(y)$ in (4) is replaced by an auxiliary variable $\mathbf{u}$.
@@
The corresponding alternating minimization steps for this Lagrangian are
[…]
@@
we propose solving (9) via Nesterov Accelerated Gradient (NAG) (Nesterov, 1983, 2007), with the details present in Algorithm 2. Of course, one could also use other optimizers such as stochastic gradient methods (Ehrhardt et al., 2025) (if the objective can be minibatched) or quasi-Newton methods (Tan et al., 2024) for this step.
However, the Lagrangian with respect to $\theta$ is nonconvex, which means the minimization over $\theta$ is often intractable. Here we bypass it by calling backpropagating over $\theta$ and taking one gradient descent step. The complete algorithm for Lagrangian-based fast equivariant imaging can be found in Algorithm 1.
Algorithm 1 Fast Equivariant Imaging (FEI)
@@
  • $\mathcal{Y}$ Observations (training dataset)
  • $\lambda$ Lagrangian multiplier
  • $\alpha$ EI regularization parameter
  • $N$ number of epochs
  • $\mathbf{u}_{0}$ initial NAG point
  • $v_{0}$ initial NAG velocity point
  • $\beta$ NAG momentum factor
  • $\eta$ NAG step size
  • $J$ number of NAG iterations
  • $\mathcal{G}$ Group of transformations
@@ 2.2 Fast Plug-and-Play Equivariant Imaging
The plug-and-play (Venkatakrishnan et al., 2013) prior is a flexible framework that incorporates state-of-the-art denoising algorithms as priors in more challenging inverse problems. Our goal here is to bring this method to Fast Equivariant Imaging, in the hope of boosting its convergence rate and its performance as well.
Starting with (5), we turn the constraint into a penalty using the Augmented Lagrangian (AL) (Afonso et al., 2010),
$\min_{\theta,\mathbf{u}}\mathbb{E}_{y\in\mathcal{Y},g\sim\mathcal{G}}\{f_{\mathrm{mc}}(y,A\mathbf{u})+\alpha\|T_{g}G_{\theta}(y)-G_{\theta}(AT_{g}G_{\theta}(y))\|_{2}^{2}+\frac{\lambda}{2}\|\mathbf{u}-G_{\theta}(y)\|_{2}^{2}+\lambda\mathbf{L}^{T}(\mathbf{u}-G_{\theta}(y))\},$ (10)
where $\mathbf{L}$ stands for the Lagrange multiplier vector for the set of equality constraints, and $\lambda$ is a free parameter to be chosen. Merging the last two terms, we get the scaled form of (10),
$\min_{\theta,\mathbf{u,L}}{\mathbb{E}_{y\in\mathcal{Y},g\sim\mathcal{G}}}\{f_{\mathrm{mc}}(y,A\mathbf{u})+\alpha\|T_{g}G_{\theta}(y)-G_{\theta}(AT_{g}G_{\theta}(y))\|_{2}^{2}+\frac{\lambda}{2}\|\mathbf{u}-G_{\theta}(y)-\mathbf{L}\|_{2}^{2}\}.$ (11)
The remedy for this problem comes in the form of the Alternating Directions Method of Multipliers (Boyd et al., 2011)(ADMM), which amounts to a sequential update of the three unknowns in this expression: $\theta$, $\mathbf{u}$, and $\mathbf{L}$. Fixing $\theta_{k},\mathbf{L}_{k}$, $\mathbf{u}$ should be updated by solving
@@
However, solving this linear system exactly often requires enormous computational efforts, which hinders the convergence of the algorithm, and we suggest solving it in one of two ways: The first option is to use the fixed-point strategy as presented in Algorithm 2, applying the Nesterov Accelerated Gradient update several times provides the necessary update for $\mathbf{u}$. An alternative approach to update $\mathbf{u}$ is Plug-and-Play, for the Lagrangian quadratic data-fidelity term (12), the classical PnP algorithm reads
@@
Fixing $\mathbf{u}_{k+1},\mathbf{L}_{k}$, the update of $\theta$ is done by solving
which can be updated by calling backpropagating and taking one gradient descent step over $\theta$.
As for the Lagrange multiplier vector $\mathbf{L}$, its update is much easier, given by
$\mathbf{L}_{k+1}=\mathbf{L}_{k}-\mathbf{u}_{k+1}+G_{\theta_{k+1}}(y)$ (16)
as emerging from the AL method (Afonso et al., 2010). Algorithm 3 summarizes the steps to be taken to apply this overall algorithm to handle minimization of the objective of equivariant imaging. Moreover, taking into account of more recent equivariant PnP (Terris et al., 2024; Tang et al., 2024), we can also modify the PnP step with equivariant denoisers. We present this option in Algorithm 4.
@@
  • $\mathcal{Y}$ Observations (training dataset)
  • $\mathbf{L}$ Augmented Lagrangian multiplier
  • $\lambda$ Augmented Lagrangian free parameter
  • $\mathcal{D}(\cdot)$ Denoiser
  • $\alpha$ EI regularization parameter
  • $N$ number of epochs
  • $\mathcal{G}$ Group of transformations
  • $\mathbf{u}$ Initial gradient descent point of (14)
  • $\gamma$ Gradient descent step size in PnP step
@@
12:    Update $\mathbf{L}_{k+1}^{(y)}=\mathbf{L}_{k}^{(y)}-x_{1}+G_{\theta_{k+1}}(y)$;
@@
  • $\mathcal{Y}$ Observations (training dataset)
  • $\mathbf{L}$ Augmented Lagrangian multiplier
  • $\lambda$ Augmented Lagrangian free parameter
  • $\mathcal{D}(\cdot)$ Denoiser
  • $\alpha$ EI regularization parameter
  • $N$ number of epochs
  • $\mathcal{G}$ Group of transformations
  • $\mathbf{u}$ Initial gradient descent point of (14)
  • $\gamma$ Gradient descent step size in PnP step
@@
12:    Update $\mathbf{L}_{k+1}^{(y)}=\mathbf{L}_{k}^{(y)}-x_{1}+G_{\theta_{k+1}}(y)$;

### v2

@@ Abstract
In this work, we propose Fast Equivariant Imaging (FEI), a novel unsupervised learning framework to rapidly and efficiently train deep imaging networks without ground-truth data. From the perspective of reformulating the Equivariant Imaging based optimization problem via the method of Lagrange multipliers and utilizing plug-and-play denoisers, this novel unsupervised scheme shows superior efficiency and performance compared to the vanilla Equivariant Imaging paradigm. In particular, our FEI schemes achieve an order-of-magnitude (10x) acceleration over standard EI on training U-Net for X-ray CT reconstruction and image inpainting, with improved generalization performance.
@@ 1 Introduction
$y\approx Ax^{\dagger},$ (1)
with reference signal / image $x^{\dagger}\in\mathbb{R}^{n}$ to be estimated, noisy measurements $y\in\mathbb{R}^{m}$. $A\in\mathbb{R}^{m\times n}$ is the forward operator potentially with $m\ll n$, and can be a downsampling operator for super-resolution task, masks for inpainting, Gaussian linear measurements for compressive sensing, Radon transform for CT/PET, to name a few. In this setting, one can construct an MAP estimator for $x^{\dagger}$ by solving the following optimization problem:
@@
where $f_{\mathrm{mc}}(Ax,y)$ is the measurement consistency (MC) term such as $\ell_{1}$ loss, $\ell_{2}$ loss and SURE-loss (Chen et al., 2022), $R(x)$ is the prior or regularizer. The paradigm presented in this paper could be easily extended to handle different types of noise (e.g., Possion, Gamma, mixed Possion-Gamma, or other models). This should be done by replacing $f_{\text{mc}}$ by the minus log-likelihood of the appropriate distribution, as done in Chen et al. (2022).
1.1 Equivariant Imaging
In unsupervised learning, since we only have access to noisy measurements $\mathcal{Y}$, one can reconstruct the signal/image by learning a reconstruction function $G_{\theta}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{m}\to\mathbb{R}^{n}$ such that $G_{\theta}(y)\approx x^{\dagger}$. By enforcing that the inverse mapping $G_{\theta}$ is consistent in the measurement domain, the following measurement consistence estimator (MC) can be obtained,
@@ 1.1 Equivariant Imaging
However, the MC constraint is not enough to learn the inverse mapping, as it cannot capture information about the signal/image outside the range of the operator $A^{\top}$, as explained in Chen et al. (2021).
In order to learn beyond the range space of $A^{\top}$, Chen et al. (2021) takes advantage of some mild prior information mainly about the symmetries of plausible signal/image, $i.e.$, signal/image is invariant to certain groups of transformations, such as shifts, rotations, reflections, etc. According to invariance, composition $G_{\theta}\circ A$ should be equivariant to the transformation $T_{g}$, $i.e.$
@@
This deep-learning-based unsupervised method has recently achieved state-of-the-art performance for a host of inverse problems (Chen et al., 2021, 2022; Terris et al., 2024). Their success has been attributed to their ability to explore group symmetry from a range of physics, at the cost of cumbersome and time-consuming training. However, one of the main challenges of this type of unsupervised schemes is the computational cost of training due to the slow convergence and large computational complexity across training iterations, particularly for large-scale inverse problems and advanced architectures such as unrolling (Monga et al., 2021; Zhou et al., 2024). In our recent work of Sketched EI (Xu et al., 2024) we partially addressed the computational complexity issue for inverse problems on which dimensionality reduction strategies can be applied. In this work, we aim to address the problem of slow convergence of standard optimizers such as Adam (Kingma and Ba, 2015) to minimize this EI objective.
1.2 Lagrange-based Splitting method
The success of EI in unsupervised learning has created a pressing need to efficiently solve (4), especially for large-scale datasets where computational scalability is critical. In this work, we explore from an optimization perspective to accelerate this promising unsupervised method via alternating minimization (Afonso et al., 2010), namely Fast Equivariant Imaging (FEI), for faster convergence.
First, it should be mentioned that the objective (4) can be seen as the Lagrangian of constrained optimization problems:
$\min_{\mathbf{u},\theta}\mathbb{E}_{y\in\mathcal{Y},g\sim\mathcal{G}}\{f_{\mathrm{mc}}(y,A\mathbf{u})+\alpha\|T_{g}G_{\theta}(y)-G_{\theta}(AT_{g}G_{\theta}(y))\|_{2}^{2}\}\quad\text{subject to}\quad\mathbf{u}=G_{\theta}(y),$ (5)
The approach proposed in this paper is based on variable splitting (Afonso et al., 2010). The idea is to split the variables $\mathbf{u}$ and $\theta$, each to serve as the argument for each of the two functions, and then minimize the sum of the two functions under the constraint that the two variables have to be equal, so that the problems are equivalent. By decoupling the measurement-consistency term as a numerical block explicitly from network complexity, enabling faster and more reliable convergence to optimal.
Our modification is achieved using an Lagrangian formulation of the problem, and is solved using an alternating-direction minimization method. From numerical experiments, we show that this fast EI method and its variant can improve both the convergence rate and the performance.
@@ 1.3 Relation to Prior Works
We differ from Chen et al. (2021) in two important ways: First, the way they minimize the objective function is by optimizing the data consistency term and the equivariance term uniformly in a single backpropagation, which hence enjoys moderate convergence and time-consuming training. In contrast, we consider a more aggressive minimization for the data consistency term via proximal alternating minimization. A second crucial difference is that within our FEI scheme, we can easily incorporate plug-and-play framework as implicit priors to further speed the convergence. In sharp contrast, the vanilla EI, by only exploiting group symmetry of the physics, can be seen as a special case of our proposed PnP-FEI. Although we focus only on accelerating EI in this work, the same algorithmic structure can be easily extended to more recent variants of EI such as Robust Equivariant Imaging (REI) (Chen et al., 2022), Multi-Operator Imaging (MOI) (Tachella et al., 2022), and our recently proposed Sketched Equivariant Imaging (SkEI) (Xu et al., 2024), which utilizes dimensionality reduction techniques for acceleration.
1.4 Contributions
We summarize three major contributions from our work:
  • We propose an efficient unsupervised training paradigm FEI which conceptually splits the original brute-force EI training into two alternating steps: a Latent-Reconstruction Step and a Pseudo-Supervision Step. We demonstrate numerically that such a splitting can massively accelerate the unsupervised training of deep imaging networks.
  • Under the FEI splitting framework, we propose two novel optimization schemes for efficiently implementing FEI, based on a new integration of adaptive gradient methods (such as Adam) with Half-Quadratic Splitting and Linearized ADMM. We utilize the gradient history of the subproblems from previous HQS/ADMM iterations to update the network parameters to ensure the fastest possible convergence.
  • Building upon FEI, we also propose the first unsupervised training scheme PnP-FEI which utilizes pretrained denoisers for further numerical improvements. Thanks to the splitting provided by FEI, we can effectively utilize pretrained image-domain priors such as deep denoiser within our framework, in the Pseudo-Supervision Step. This is the first paradigm which utilizes both the primal domain (image) prior and the dual domain (measurement) prior for unsupervised training of deep imaging networks.
@@ 2 Proposed Methods
In this section, we present our FEI framework which is based on splitting the original objective of EI into two alternating subproblems: Latent-reconstruction and Pseudo-supervision. This leads to two options for FEI algorithmically, one with a style of Half-Quaratic Splitting, and one with a style of Linearized ADMM.
2.1 Fast Equivariant Imaging — Option 1 ( Adaptive Half Quadratic Splitting)
We first reformulate (4) by substituting $\mathbf{u}=G_{\theta}(y)$ as follows
@@ 2.1 Fast Equivariant Imaging — Option 1 ( Adaptive Half Quadratic Splitting)
Typical choices for the measurement consistency term $f_{\mathrm{mc}}(x)$ include $\ell_{1}$ loss, $\ell_{2}$ loss or SURE-loss (Chen et al., 2022), etc. We simply substitute $G_{\theta}(y)$ with $\mathbf{u}$ in the MC term and further incorporate the EI regularizer. From another perspective, the neural network $G_{\theta}(y)$ in (4) is replaced by an auxiliary variable $\mathbf{u}$.
@@
We then address problem (7) with half quadratic splitting (HQS) algorithm due to its simplicity. Note that a very large $\lambda$ will force $\mathbf{u}$ to be approximately equal to $G_{\theta}(y)$. Usually, $\lambda$ varies in a non-descending order during the following iterative solution
$\mathbf{u}_{t+1}$ $\approx\arg\min_{\mathbf{u}}\mathcal{L}_{\lambda,\alpha}(\mathbf{u},\theta_{t});\ \ \rightarrow\mathrm{\it\color[rgb]{.75,0,.25}Latent\ reconstruction\ step}$ (8)
$\theta_{t+1}$ $\approx\arg\min_{\theta}\mathcal{L}_{\lambda,\alpha}(\mathbf{u}_{t+1},\theta);\ \ \rightarrow\mathrm{\it\color[rgb]{.75,0,.25}Pseudo\ supervision\ step}$ (9)
@@
we propose solving (10) via Nesterov Accelerated Gradient (NAG) (Nesterov, 1983, 2007), with the details present in Algorithm 2. Of course, one could also use other optimizers, such as stochastic gradient methods (Ehrhardt et al., 2025) (if the objective can be minibatched) or quasi-Newton methods (Tan et al., 2024) for this step.
However, the Lagrangian with respect to $\theta$ is nonconvex, which means the minimization over $\theta$ is often intractable. Here we bypass it by calling backpropagating over $\theta$ and taking one gradient descent step using an adaptive gradient method (Adam by default). The complete algorithm for Lagrangian-based fast equivariant imaging can be found in Algorithm 1. Note that in each iteration we typically just run 1 iteration of Adam for the Pseudo-supervision step, while the gradient and momentum history of Adam is inherited from previous Pseudo-supervision steps.
Algorithm 1 Fast Equivariant Imaging (FEI-Option 1)
@@
  • $\mathcal{Y}$ - Observations (training dataset)
  • $\lambda$ - Lagrangian multiplier
  • $\alpha$ - EI regularization parameter
  • NN - number of epochs
  • $\mathbf{u}_{0}$ - initial NAG point
  • $v_{0}$ - initial NAG velocity point
  • $\beta$ - NAG momentum factor
  • $\eta$ - NAG step size
  • $J$ - number of NAG iterations
  • $\mathcal{G}$ - Group of transformations
@@
2.2 Fast Equivariant Imaging — Option 2 ( Adaptive Linearized ADMM Splitting)
First, starting with (6), we turn the constraint into a penalty using the Augmented Lagrangian (AL) (Afonso et al., 2010),
$\mathcal{L}_{\lambda}(\mathbf{u},\theta,\mathbf{L})\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{y\in\mathcal{Y},g\in\mathcal{G}}\{f_{\textbf{mc}}(y,A\mathbf{u})+\alpha\|T_{g}G_{\theta}(y)-G_{\theta}(AT_{g}G_{\theta}(y))\|_{2}^{2}+\langle\mathbf{L},\mathbf{u}-G_{\theta}(y)\rangle+\frac{\lambda}{2}\|\mathbf{u}-G_{\theta}(y)\|_{2}^{2}$ (11)
for a penalty weight $\lambda>0$. Applied to (11), every iteration of ADMM would minimize the Augmented Lagrange with respect to $\mathbf{u}$, then with respect to $\theta$, and then update the dual variable $\mathbf{L}$, which are presented as follows:
$\mathbf{u}_{k+1}$ $\in\mathop{\arg\min}_{\mathbf{u}}\mathcal{L}_{\lambda}(\mathbf{u},\theta_{k},\mathbf{L}_{k}),\ \ \rightarrow\mathrm{\it\color[rgb]{.75,0,.25}Latent\ reconstruction\ step}$ (12)
$\theta_{k+1}$ $\in\mathop{\arg\min}_{\theta}\mathcal{L}_{\lambda}(\mathbf{u}_{k+1},\theta,\mathbf{L}_{k}),\ \ \rightarrow\mathrm{\it\color[rgb]{.75,0,.25}Pseudo\ supervision\ step}$
$\mathbf{L}_{k+1}$ $=\mathbf{L}_{k}-\tau(\mathbf{u}_{k+1}-G_{\theta_{k+1}}(y)).$
then take only one descent step in both $\mathbf{u}$ and $\theta$. Specifically, fixing $\theta_{k},\mathbf{L}_{k}$, the update of $\mathbf{u}_{k+1}$ can be performed by one step gradient descent, then fixing $\mathbf{u}_{k+1},\mathbf{L}_{k}$, the update of $\theta$ is done by solving
which can be updated by calling backpropagating and taking one adaptive gradient descent (Adam by default) step $\theta$. Note that in each iteration we typically just run 1 iteration of Adam for the Pseudo-supervision step, while the gradient and momentum history of Adam is inherited from previous Pseudo-supervision steps. As for the Lagrange multiplier vector $\mathbf{L}$, its update is much easier, given by
$\mathbf{L}_{k+1}=\mathbf{L}_{k}-\tau(\mathbf{u}_{k+1}-G_{\theta_{k+1}}(y))$ (14)
as emerging from the AL method (Afonso et al., 2010).
Algorithm 3 Fast Equivariant Imaging (FEI-Option 2)
1:Inputs:
  • $\mathcal{Y}$ - Observations (training dataset)
  • $\mathbf{L}$ - Augmented Lagrangian multiplier
  • $\lambda$ - Augmented Lagrangian free parameter
  • $\alpha$ - EI regularization parameter
  • NN - number of epochs
  • $\mathcal{G}$ - Group of transformations
  • $\mathbf{u}$ - Initial gradient descent point of (18)
  • $\gamma$ - Gradient descent step size
3:For $i=0,\cdots,N-1$ do:
4:  For $y\in\mathcal{Y}$ do:
5:    Sample $T_{g}$ where $g\sim\mathcal{G}$;
6:    Compute $x_{0}=G_{\theta_{k}}(y)$, by default let $\mathbf{u}_{k}=x_{0}$;
7:    Update $\mathbf{u}_{k+1}=\mathbf{u}_{k}-\gamma[\nabla f_{\mathrm{mc}}(y,A\mathbf{u}_{k})+\lambda(\mathbf{u}_{k}-G_{\theta_{k}}(y)-\mathbf{L}_{k}^{(y)})]$
8:    Let $x_{1}=\mathbf{u}_{k+1}$;
9:    Transform $x_{2}=T_{g}(x_{1})$;
10:    Compute $x_{3}=G_{\theta_{k}}(Ax_{2})$;
12:    Update $\mathbf{L}_{k+1}^{(y)}=\mathbf{L}_{k}^{(y)}-\tau(x_{1}-G_{\theta_{k+1}}(y))$;
13:    $k\leftarrow k+1$
14:  End For
15:End For
16:Output:  $G_{\theta^{\star}}$
@@ 2.3 Fast Plug-and-Play Equivariant Imaging
The plug-and-play (Venkatakrishnan et al., 2013) prior is a flexible framework that incorporates state-of-the-art denoising algorithms as priors in more challenging inverse problems. Our goal here is to bring this method to Fast Equivariant Imaging, in the hope of boosting its convergence rate and its performance as well. Using the last two terms of (11), we get the scaled form,
$\min_{\theta,\mathbf{u}}{\mathbb{E}_{y\in\mathcal{Y},g\sim\mathcal{G}}}\{f_{\mathrm{mc}}(y,A\mathbf{u})+\alpha\|T_{g}G_{\theta}(y)-G_{\theta}(AT_{g}G_{\theta}(y))\|_{2}^{2}+\frac{\lambda}{2}\|\mathbf{u}-G_{\theta}(y)-\mathbf{L}\|_{2}^{2}\}.$ (15)
The remedy for this problem comes in the form of ADMM, which amounts to a sequential update of the three unknowns in this expression: $\theta$, $\mathbf{u}$, and $\mathbf{L}$. Fixing $\theta_{k},\mathbf{L}_{k}$, $\mathbf{u}$ should be updated by solving
@@
However, solving this linear system exactly often requires enormous computational efforts, which hinders the convergence of the algorithm, and we suggest solving it in one of two ways: The first option is to use the one gradient-descent step strategy as presented in Algorithm 3. An alternative approach to update $\mathbf{u}$ is Plug-and-Play, for the Lagrangian quadratic data-fidelity term (16), the classical PnP algorithm reads
@@
The Algorithm 4 summarizes the steps to be taken to apply this overall algorithm to manage the minimization of the objective of equivariant imaging. Moreover, taking into account the more recent equivariant PnP (Terris et al., 2024; Tang et al., 2024), we can also modify the PnP step with equivariant denoisers. We present this option in Algorithm 5.
@@
  • $\mathcal{Y}$ - Observations (training dataset)
  • $\mathbf{L}$ - Augmented Lagrangian multiplier
  • $\lambda$ - Augmented Lagrangian free parameter
  • $\mathcal{D}(\cdot)$ - Denoiser
  • $\alpha$ - EI regularization parameter
  • NN - number of epochs
  • $\mathcal{G}$ - Group of transformations
  • $\mathbf{u}$ - Initial gradient descent point of (18)
  • $\gamma$ - Gradient descent step size in PnP step
@@
12:    Update $\mathbf{L}_{k+1}^{(y)}=\mathbf{L}_{k}^{(y)}-\tau(x_{1}-G_{\theta_{k+1}}(y))$;
@@
  • $\mathcal{Y}$ - Observations (training dataset)
  • $\mathbf{L}$ - Augmented Lagrangian multiplier
  • $\lambda$ - Augmented Lagrangian free parameter
  • $\mathcal{D}(\cdot)$ - Denoiser
  • $\alpha$ - EI regularization parameter
  • NN - number of epochs
  • $\mathcal{G}$ - Group of transformations
  • $\mathbf{u}$ - Initial gradient descent point of (18)
  • $\gamma$ - Gradient descent step size in PnP step
@@
12:    Update $\mathbf{L}_{k+1}^{(y)}=\mathbf{L}_{k}^{(y)}-\tau(x_{1}-G_{\theta_{k+1}}(y))$;